William​ Beville's computer training​ school, in​ Richmond, stocks workbooks with the following​ characteristics:

Demand D

19,400

​units/year

Ordering cost S

​$25​/order

Holding cost H

​$4​/unit/year

​a) The EOQ for the workbooks is ___ ​(round your response to the nearest whole​ number).​

b) What are the annual holding costs for the​ workbooks? ​$___ ​(round your response to the nearest whole​ number).​

c) What are the annual ordering​ costs? ​$___ (round your response to the nearest whole​ number).
Answer #1

Given,

Demand, D = 19400 units/year

Ordering cost, P = $25/order

Holding cost, C = $4/unit per year

==========================================================================

a) EOQ

EOQ = V 2DP с • D = Demand in units for specified period • P=Relevant ordering costs per PO • C= Relevant carrying costs of o

Hence, substituting the above values in the formula,

EOQ = SqRoot((2(19400)(25)/4)= SqRoot(242500)

Therefore EOQ = 492.44289008980525 = 492

======================================================================

b) Annual Holding Cost :

Q Annual Holding Cost H 2

= 492 /2 * 4

Therefore , Annual Holding Cost = $ 984

==================================================================

c) Annual Ordering Cost:

D Annual Ordering Cost x S Q

= 19400/492 *25

Therefore, Annual Ordering Cost = 985.7723577235772 = $ 986

===================================================================================

I hope it helps.

Please UPVOTE as a token of appreciation.

Thank you.

Good day !